# ProtoTS: Learning Hierarchical Prototypes for Explainable Time Series Forecasting

### v3

@@ Abstract
While deep learning has achieved impressive performance in time series forecasting, it becomes increasingly crucial to understand its decision-making process for building trust in high-stakes scenarios. Existing interpretable models often provide only local and partial explanations, lacking the capability to reveal how heterogeneous and interacting input variables jointly shape the overall temporal patterns in the forecast curve. We propose ProtoTS, a novel interpretable forecasting framework that achieves both high accuracy and transparent decision-making through modeling prototypical temporal patterns. ProtoTS computes instance-prototype similarity based on a denoised representation that preserves abundant heterogeneous information. The prototypes are organized hierarchically to capture global temporal patterns with coarse prototypes while capturing finer-grained local variations with detailed prototypes, enabling expert steering and multi-level interpretability. Experiments on multiple realistic benchmarks, including a newly released LOF dataset, show that ProtoTS not only exceeds existing methods in forecast accuracy but also delivers expert-steerable interpretations for better model understanding and decision support.
@@ 1 Introduction
Time series forecasting has been widely applied in high-stakes scenarios such as load forecasting (jiang2024interpretable; yang2023interactive), energy management (deb2017review; weron2014electricity), weather prediction (angryk2020multivariate; karevan2020transductive), all of which involve considerable financial impacts. In these applications, while achieving high forecast accuracy is crucial, understanding why and how the model makes specific predictions is equally important. It aids in preventing substantial financial losses and building the trust necessary (rojat2021explainable).
A range of explainable time series forecasting methods have been developed to simultaneously ensure interpretability and good predictive performance (oreshkin2019n; lim2021temporal; zhao2024disentangled; lin2024cyclenet). However, their overall interpretability and potential for further performance improvement are limited, since they mainly provide local, partial explanations for both the output and input sides:
  • C1: For the output side, existing methods (lim2021temporal; zhao2024disentangled) mainly explain the prediction at individual time steps, lacking the ability to help users quickly interpret the reasons behind the overall trend in the forecast curve. For example, they fail to explain why three peaks occur in the predicted electricity load curve at the noon, afternoon, and night, respectively, and why the three peaks gradually descend (Figure 1(a1)). Understanding such overall temporal patterns is of great importance in real-world scenarios. For example, power system experts need to ensure demand peaks are identified correctly, so that they can make dispatch decisions (6721569), such as whether additional electricity should be purchased from external sources. Failing to provide such analyses limits interpretability and prevents experts from steering the model to improve its accuracy.
  • C2: For the input side, existing explanations are often limited to certain types of variables, such as focusing solely on endogenous variables (lin2024cyclenet). However, as shown in Figure 1(a), there are many different types of input variables, and understanding how their interactions impact the temporal patterns is critical. For example, the typical temporal pattern of extreme summer conditions can only be observed when both temperature and seasonal variables are considered together.
@@
To address these challenges, we propose a Prototypical Time Series Forecasting framework (ProtoTS) that effectively models the overall temporal patterns (C1) and enables prototypes to capture the interactions among inputs (C2). As shown in Figure 1(b), each of our prototypes corresponds to a typical temporal pattern. For example, the curve of the Spring Festival prototype [^1] The name “Spring Festival” is now manually summarized based on relevant covariates and temporal patterns, though it can also be generated using large language models. indicates that there are peaks in the morning and evening during the Spring Festival period. This pattern occurs at specific combinations of covariates (i.e., when the inputs indicate both a holiday and the first month of the lunar calendar), enabling it to explain both the formation of the overall temporal pattern and how it is influenced by the combination of multiple input features (C2). These few prototypes provide a global explanation by summarizing typical patterns across the entire dataset, thereby supporting expert intervention (e.g., refining the temporal pattern of the prototype, Sec. 4.4). At the same time, the model improves accuracy by effectively modeling the local interactions of covariates.
@@ 2 Related Work
Time Series Forecasting with Exogenous Variables Time series forecasting with exogenous variables has been extensively explored in both classical and modern approaches. Traditional statistical models such as ARIMAX (williams2001multivariate) and SARIMAX (vagropoulos2016comparison) extend the ARIMA framework to incorporate correlations between exogenous and endogenous variables. Seeking richer dynamics, Transformer families like TFT (lim2021temporal) and TimeXer (wang2024timexer), introduce attention mechanisms to tackle this task. Parallel work in MLPs pursues efficiency and long-horizon stability: NBEATSx (oreshkin2019n) appends a dedicated stack for auxiliary factors on top of a residual basis expansion, TiDE (das2023long) factorizes temporal and feature projections with two dense encoders to harness known‑future variables, and TSMixer (ekambaram2023tsmixer; chen2023tsmixer) alternates mixing layers to effectively model their interactions. Outside deep learning, generic tabular learners such as XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), and GAM (yang2023interactive), perform well when exogenous variables are informative and strongly correlated with the target. Our method, ProtoTS, achieves strong performance on this task while offering interpretability and steerability for domain experts, effectively handling large amounts of heterogeneous variables while filtering out irrelevant information.
@@
Time Series Interpretability Time series interpretability can be divided into post-hoc and ante-hoc approaches. Post-hoc interpretability focuses on unveiling the reasoning of pre-existing black-box models (gu2025explainable). Generic explanation techniques such as SHAP (lundberg2017unified), LIME (ribeiro2016should), IG (sundararajan2017axiomatic) can be directly applied to time-series models. crabbe2021explaining learns sparse perturbation masks over the time-feature dimensions, struggling to quantify each input’s effect on a regression forecast. Consequently, regression tasks tend to favour ante-hoc models whose interpretability is built in by design. Following the attribution-centric view of post‑hoc explanations, TFT (lim2021temporal) provides explorable attention over input time steps, and DiPE-Linear (zhao2024disentangled) parameterizes filters in both temporal and spectral domains to visualise the regions of interest. N-BEATS (oreshkin2019n) uses residual blocks for generic, trend, and seasonal components, and olivares2023neural extends the scheme to exogenous variables. These works are limited to local explanations and fail to provide a global view that experts need for a shared understanding. CycleNet (lin2024cyclenet) identifies common periodic patterns in time-series data but only focuses on endogenous variables. In contrast, ProtoTS holds a global set of prototypical patterns and explains each forecast by matching it to them, offering both global overview and local interpretation. Its hierarchical structure supports expert steering at multiple levels, ensuring faithful overall understanding.
@@
Prototypical Time Series Model Previous works have already shown the value of prototypes in time series models. These methods typically fall into two categories. 1) prototypes decoded as intermediate representations (shen2025learn; li2023prototype; jin2023time) or model parameters (chen2024similarity). For example, jin2023time uses prototypes to map time-series embeddings into textual embeddings for large language models. Such designs enrich the model’s architecture but leave the prototypes detached from the output. 2) prototypes directly mapped to output variables (queen2023encoding; liu2024timex++; ming2019interpretable; obermair2023example), serving as typical examples of the outputs. While the latter offers stronger interpretability, existing approaches decode a prototype into only a single output variable (e.g., a class prediction). In contrast, our method is, to the best of our knowledge, the first to decode a prototype into a sequence of output variables (e.g., a 96-step forecasting curve), enabling holistic interpretation of temporal dynamics.
@@ 3.1 Formulation and Framework
Problem Formulation We consider the task of time series forecasting with exogenous variables, as shown in Figure 1(a), since abundant heterogeneous exogenous variables provide rich information to boost endogenous predictability. Formally, let the historical endogenous variables (e.g., electric load) as $\mathbf{Y}_{1:L}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\}\in\mathbb{R}^{L\times 1}$ and the associated exogenous variables (e.g., temperature and holiday) as $\mathbf{X}_{1:L}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{L}\}\in\mathbb{R}^{L\times C}$, where $L$ is the look-back window size and $C$ is the number of exogenous variables. In many practical scenarios (wessel2020using; lin2023prediction), future exogenous variables $\mathbf{X}_{L+1:L+H}=\{\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+H}\}\in\mathbb{R}^{H\times C}$ are also available. For example, in the LOF dataset, we obtain forecasted weather variables (e.g., temperature) from commercial weather services, and calendar variables (e.g., holiday flags) are accessible in advance. In our setting, we explicitly incorporate future exogenous variables, as these substantially improve forecasting performance. The goal is to predict the future curve of the endogenous variable for the next $H$ time steps, denoted as $\mathbf{Y}_{L+1:L+H}=\{\mathbf{y}_{L+1},\dots,\mathbf{y}_{L+H}\}\in\mathbb{R}^{H\times 1}$. The forecasting model $\mathcal{F}$ takes the input variables to predict the future endogenous variables $\widehat{\mathbf{Y}}$:
@@ 3.2 Multi-Channel Prototype Similarity Computation
Multi-Channel Embedding In ProtoTS, we adopt a multi-channel embedding where endogenous and exogenous variables are processed separately to achieve optimal encoding for each type. Specifically, endogenous variables are directly encoded through an endogenous encoding channel, implemented as a non-linear projection $\gamma$ using a multi-layer perceptron with activation functions, a widely used design (das2023long). For exogenous variables, we follow standard practice in tabular learning (wang2022transtab). We categorize them into discrete variables (e.g., is holiday, day of week) and continuous variables (e.g., temperature, humidity), and apply tailored embedding methods to best capture the properties of each variable. Concretely, exogenous variables $\mathbf{x}_{t}$ at time $t$ are decomposed into $\mathbf{x}_{t}^{\text{dis}}\in\mathbb{N}^{C_{\text{dis}}}$ and $\mathbf{x}^{\text{con}}_{t}\in\mathbb{R}^{C_{\text{con}}}$, $C_{\text{dis}}$ and $C_{\text{con}}$ refer to the number of discrete and continuous variables. Discrete variable $\mathbf{x}_{t,j}^{\text{dis}}$ are processed through a discrete encoding channel using dedicated embedding tables $\mathbf{E}_{j}\in\mathbb{R}^{|\Omega_{j}|\times d}$, where $\Omega_{j}$ is the vocabulary and $d$ is the embedding dimension. Continuous variables $\mathbf{x}^{\text{con}}_{t,j}$ are projected into the embedding space through a continuous encoding channel, applying variable-specific non-linear projections $\psi_{j}$. This multi-channel embedding allows the model to better preserve the information contained in heterogeneous variables.
Bottleneck Channel Fusion To effectively fuse all variables information at time step $t$ and obtain a full representation, we aggregate the embeddings of all variables through addition, which integrates information across heterogeneous variables without additional parameters (arora2017simple):
@@
During the forecast window, where $\mathbf{y}_{t}$ is the target to be predicted, only exogenous variables are used to construct $\mathbf{Z}_{t}$. The embedding $\mathbf{Z}_{1:L+H}=\{\mathbf{Z}_{1},\cdots,\mathbf{Z}_{t},\cdots,\mathbf{Z}_{L+H}\}\in\mathbb{R}^{(L+H)\times d}$ now encodes a rich amount of information from endogenous and exogenous variables. However, such a rich embedding may also introduce noise, as irrelevant information from heterogeneous inputs can be mixed into the representations. To mitigate this, ProtoTS introduces a bottleneck layer to filter out irrelevant information while retaining the most predictive components. Concretely, we project the high-dimensional embeddings into a lower-dimensional latent space through a bottleneck projection, $\mathbb{R}^{d}\to\mathbb{R}^{d_{\text{bottle}}}\to\mathbb{R}^{d}$, where $d_{\text{bottle}}\ll d$. The compressed representation is then mapped back to the original space, retaining the most relevant information while filtering out noise (wang2021variational). We apply the bottleneck layer within an MLP-Mixer architecture (tolstikhin2021mlp), inserting it into both $\text{MLP}_{\text{feature}}$ and $\text{MLP}_{\text{time}}$. Fusion is performed sequentially, first along the feature dimension and then along the temporal dimension, T denotes the transpose of the vector:
@@
Prototype Similarity Computing To provide a user-friendly global explanation while handling complex information at a local level for detailed interpretation and accurate forecasting, ProtoTS integrates prototypical explanation directly into its prediction mechanism via a learnable set of prototypes $\mathbf{\Pi}$. Each prototype in $\mathbf{\Pi}$ is represented as an embedding $\bm{\mu}\in\mathbb{R}^{d}$ and its corresponding temporal pattern $\mathbf{p}\in\mathbb{R}^{T}$, which are both learnable parameters. Here, $\bm{\mu}$ is used to compute the similarity between the query representation $\hat{\mathbf{Z}}$ and the prototypes, by applying a softmax over their distances $d(\hat{\mathbf{Z}},\bm{\mu})$. The distance can be chosen flexibly, such as Euclidean distance, cosine similarity. We adopt Euclidean distance, under which the prototypes are equivalent to a Gaussian mixture model in the embedding with an identity covariance matrix (allen2019infinite). Meanwhile, temporal pattern $\mathbf{p}$ is used to form the final prediction. Specifically, we fix each prototype temporal pattern $\mathbf{p}$ over a predefined period of length $T$, and during forecasting, the prototype pattern is aligned with the forecast window $H$ based on its phase and length (lin2024cyclenet).

### v4

# arXiv:2509.23159v4[cs.LG] 27 Feb 2026

@@ Abstract
While deep learning has achieved impressive performance in time series forecasting, it becomes increasingly crucial to understand its decision-making process for building trust in high-stakes scenarios. Existing interpretable models often provide only local and partial explanations, lacking the capability to reveal how heterogeneous and interacting input variables jointly shape the overall temporal patterns in the forecast curve. We propose ProtoTS, a novel interpretable forecasting framework that achieves both high accuracy and transparent decision-making through modeling prototypical temporal patterns. ProtoTS computes instance-prototype similarity based on a denoised representation that preserves abundant heterogeneous information. The prototypes are organized hierarchically to capture global temporal patterns with coarse prototypes while capturing finer-grained local variations with detailed prototypes, enabling expert steering and multi-level interpretability. Experiments on multiple realistic benchmarks, including a newly released LOF dataset, show that ProtoTS not only exceeds existing methods in forecast accuracy but also delivers expert-steerable interpretations for better model understanding and decision support. The source code is available at https://github.com/SKURA502/ProtoTS.
@@ 1 Introduction
Time series forecasting has been widely applied in high-stakes scenarios such as load forecasting (Jiang et al., 2024; Yang et al., 2023), energy management (Deb et al., 2017; Weron, 2014), weather prediction (Angryk et al., 2020; Karevan and Suykens, 2020), all of which involve considerable financial impacts. In these applications, while achieving high forecast accuracy is crucial, understanding why and how the model makes specific predictions is equally important. It aids in preventing substantial financial losses and building the trust necessary (Rojat et al., 2021).
A range of explainable time series forecasting methods have been developed to simultaneously ensure interpretability and good predictive performance (Oreshkin et al., 2019; Lim et al., 2021; Zhao et al., 2024; Lin et al., 2024). However, their overall interpretability and potential for further performance improvement are limited, since they mainly provide local, partial explanations for both the output and input sides:
  • C1: For the output side, existing methods (Lim et al., 2021; Zhao et al., 2024) mainly explain the prediction at individual time steps, lacking the ability to help users quickly interpret the reasons behind the overall trend in the forecast curve. For example, they fail to explain why three peaks occur in the predicted electricity load curve at the noon, afternoon, and night, respectively, and why the three peaks gradually descend (Figure 1(a1)). Understanding such overall temporal patterns is of great importance in real-world scenarios. For example, power system experts need to ensure demand peaks are identified correctly, so that they can make dispatch decisions (Thanos et al., 2013), such as whether additional electricity should be purchased from external sources. Failing to provide such analyses limits interpretability and prevents experts from steering the model to improve its accuracy.
  • C2: For the input side, existing explanations are often limited to certain types of variables, such as focusing solely on endogenous variables (Lin et al., 2024). However, as shown in Figure 1(a), there are many different types of input variables, and understanding how their interactions impact the temporal patterns is critical. For example, the typical temporal pattern of extreme summer conditions can only be observed when both temperature and seasonal variables are considered together.
@@
To address these challenges, we propose a Prototypical Time Series Forecasting framework (ProtoTS) that effectively models the overall temporal patterns (C1) and enables prototypes to capture the interactions among inputs (C2).
As shown in Figure 1(b), each of our prototypes corresponds to a typical temporal pattern. For example, the curve of the Spring Festival prototype [^1] The name “Spring Festival” is now manually summarized based on relevant covariates and temporal patterns, though it can also be generated using large language models. indicates that there are peaks in the morning and evening during the Spring Festival period. This pattern occurs at specific combinations of covariates (i.e., when the inputs indicate both a holiday and the first month of the lunar calendar), enabling it to explain both the formation of the overall temporal pattern and how it is influenced by the combination of multiple input features (C2). These few prototypes provide a global explanation by summarizing typical patterns across the entire dataset, thereby supporting expert intervention (e.g., refining the temporal pattern of the prototype, Sec. 4.4). At the same time, the model improves accuracy by effectively modeling the local interactions of covariates.
@@ 2 Related Work
Time Series Forecasting with Exogenous Variables Time series forecasting with exogenous variables has been extensively explored in both classical and modern approaches. Traditional statistical models such as ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) extend the ARIMA framework to incorporate correlations between exogenous and endogenous variables. Seeking richer dynamics, Transformer families like TFT (Lim et al., 2021) and TimeXer (Wang et al., 2024), introduce attention mechanisms to tackle this task. Parallel work in MLPs pursues efficiency and long-horizon stability: NBEATSx (Oreshkin et al., 2019) appends a dedicated stack for auxiliary factors on top of a residual basis expansion, TiDE (Das et al., 2023) factorizes temporal and feature projections with two dense encoders to harness known‑future variables, and TSMixer (Ekambaram et al., 2023; Chen et al., 2023) alternates mixing layers to effectively model their interactions. Outside deep learning, generic tabular learners such as XGBoost (Chen and Guestrin, 2016), LightGBM (Ke et al., 2017), and GAM (Yang et al., 2023), perform well when exogenous variables are informative and strongly correlated with the target. Our method, ProtoTS, achieves strong performance on this task while offering interpretability and steerability for domain experts, effectively handling large amounts of heterogeneous variables while filtering out irrelevant information.
@@
Time Series Interpretability Time series interpretability can be divided into post-hoc and ante-hoc approaches. Post-hoc interpretability focuses on unveiling the reasoning of pre-existing black-box models (Gu et al., 2025). Generic explanation techniques such as SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), IG (Sundararajan et al., 2017) can be directly applied to time-series models. Crabbé and Van Der Schaar (2021) learns sparse perturbation masks over the time-feature dimensions, struggling to quantify each input’s effect on a regression forecast. Consequently, regression tasks tend to favour ante-hoc models whose interpretability is built in by design. Following the attribution-centric view of post‑hoc explanations, TFT (Lim et al., 2021) provides explorable attention over input time steps, and DiPE-Linear (Zhao et al., 2024) parameterizes filters in both temporal and spectral domains to visualise the regions of interest. N-BEATS (Oreshkin et al., 2019) uses residual blocks for generic, trend, and seasonal components, and Olivares et al. (2023) extends the scheme to exogenous variables. These works are limited to local explanations and fail to provide a global view that experts need for a shared understanding. CycleNet (Lin et al., 2024) identifies common periodic patterns in time-series data but only focuses on endogenous variables. In contrast, ProtoTS holds a global set of prototypical patterns and explains each forecast by matching it to them, offering both global overview and local interpretation. Its hierarchical structure supports expert steering at multiple levels, ensuring faithful overall understanding.
@@
Prototypical Time Series Model Previous works have already shown the value of prototypes in time series models. These methods typically fall into two categories. 1) prototypes decoded as intermediate representations (Shen, 2025; Li et al., 2023a; Jin et al., 2023a) or model parameters (Chen et al., 2024). For example, Jin et al. (2023a) uses prototypes to map time-series embeddings into textual embeddings for large language models. Such designs enrich the model’s architecture but leave the prototypes detached from the output. 2) prototypes directly mapped to output variables (Queen et al., 2023; Liu et al., 2024; Ming et al., 2019; Obermair et al., 2023), serving as typical examples of the outputs. While the latter offers stronger interpretability, existing approaches decode a prototype into only a single output variable (e.g., a class prediction). In contrast, our method is, to the best of our knowledge, the first to decode a prototype into a sequence of output variables (e.g., a 96-step forecasting curve), enabling holistic interpretation of temporal dynamics.
@@ 3.1 Formulation and Framework
Problem Formulation We consider the task of time series forecasting with exogenous variables, as shown in Figure 1(a), since abundant heterogeneous exogenous variables provide rich information to boost endogenous predictability. Formally, let the historical endogenous variables (e.g., electric load) as $\mathbf{Y}_{1:L}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\}\in\mathbb{R}^{L\times 1}$ and the associated exogenous variables (e.g., temperature and holiday) as $\mathbf{X}_{1:L}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{L}\}\in\mathbb{R}^{L\times C}$, where $L$ is the look-back window size and $C$ is the number of exogenous variables. In many practical scenarios (Wessel, 2020; Lin et al., 2023), future exogenous variables $\mathbf{X}_{L+1:L+H}=\{\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+H}\}\in\mathbb{R}^{H\times C}$ are also available. For example, in the LOF dataset, we obtain forecasted weather variables (e.g., temperature) from commercial weather services, and calendar variables (e.g., holiday flags) are accessible in advance. In our setting, we explicitly incorporate future exogenous variables, as these substantially improve forecasting performance. The goal is to predict the future curve of the endogenous variable for the next $H$ time steps, denoted as $\mathbf{Y}_{L+1:L+H}=\{\mathbf{y}_{L+1},\dots,\mathbf{y}_{L+H}\}\in\mathbb{R}^{H\times 1}$. The forecasting model $\mathcal{F}$ takes the input variables to predict the future endogenous variables $\widehat{\mathbf{Y}}$:
@@ 3.2 Multi-Channel Prototype Similarity Computation
Multi-Channel Embedding In ProtoTS, we adopt a multi-channel embedding where endogenous and exogenous variables are processed separately to achieve optimal encoding for each type. Specifically, endogenous variables are directly encoded through an endogenous encoding channel, implemented as a non-linear projection $\gamma$ using a multi-layer perceptron with activation functions, a widely used design (Das et al., 2023). For exogenous variables, we follow standard practice in tabular learning (Wang and Sun, 2022). We categorize them into discrete variables (e.g., is holiday, day of week) and continuous variables (e.g., temperature, humidity), and apply tailored embedding methods to best capture the properties of each variable. Concretely, exogenous variables $\mathbf{x}_{t}$ at time $t$ are decomposed into $\mathbf{x}_{t}^{\text{dis}}\in\mathbb{N}^{C_{\text{dis}}}$ and $\mathbf{x}^{\text{con}}_{t}\in\mathbb{R}^{C_{\text{con}}}$, $C_{\text{dis}}$ and $C_{\text{con}}$ refer to the number of discrete and continuous variables. Discrete variable $\mathbf{x}_{t,j}^{\text{dis}}$ are processed through a discrete encoding channel using dedicated embedding tables $\mathbf{E}_{j}\in\mathbb{R}^{|\Omega_{j}|\times d}$, where $\Omega_{j}$ is the vocabulary and $d$ is the embedding dimension. Continuous variables $\mathbf{x}^{\text{con}}_{t,j}$ are projected into the embedding space through a continuous encoding channel, applying variable-specific non-linear projections $\psi_{j}$. This multi-channel embedding allows the model to better preserve the information contained in heterogeneous variables Munz-Körner and Weiskopf (2024).
Bottleneck Channel Fusion To effectively fuse all variables information at time step $t$ and obtain a full representation, we aggregate the embeddings of all variables through addition, which integrates information across heterogeneous variables without additional parameters (Arora et al., 2017):
@@
During the forecast window, where $\mathbf{y}_{t}$ is the target to be predicted, only exogenous variables are used to construct $\mathbf{Z}_{t}$. The embedding $\mathbf{Z}_{1:L+H}=\{\mathbf{Z}_{1},\cdots,\mathbf{Z}_{t},\cdots,\mathbf{Z}_{L+H}\}\in\mathbb{R}^{(L+H)\times d}$ now encodes a rich amount of information from endogenous and exogenous variables. However, such a rich embedding may also introduce noise, as irrelevant information from heterogeneous inputs can be mixed into the representations. To mitigate this, ProtoTS introduces a bottleneck layer to filter out irrelevant information while retaining the most predictive components. Concretely, we project the high-dimensional embeddings into a lower-dimensional latent space through a bottleneck projection, $\mathbb{R}^{d}\to\mathbb{R}^{d_{\text{bottle}}}\to\mathbb{R}^{d}$, where $d_{\text{bottle}}\ll d$. The compressed representation is then mapped back to the original space, retaining the most relevant information while filtering out noise (Wang et al., 2021). We apply the bottleneck layer within an MLP-Mixer architecture (Tolstikhin et al., 2021), inserting it into both $\text{MLP}_{\text{feature}}$ and $\text{MLP}_{\text{time}}$. Fusion is performed sequentially, first along the feature dimension and then along the temporal dimension, T denotes the transpose of the vector:
@@
Prototype Similarity Computing To provide a user-friendly global explanation while handling complex information at a local level for detailed interpretation and accurate forecasting, ProtoTS integrates prototypical explanation directly into its prediction mechanism via a learnable set of prototypes $\mathbf{\Pi}$ (Zhang et al., 2024b). Each prototype in $\mathbf{\Pi}$ is represented as an embedding $\bm{\mu}\in\mathbb{R}^{d}$ and its corresponding temporal pattern $\mathbf{p}\in\mathbb{R}^{T}$, which are both learnable parameters. Here, $\bm{\mu}$ is used to compute the similarity between the query representation $\hat{\mathbf{Z}}$ and the prototypes, by applying a softmax over their distances $d(\hat{\mathbf{Z}},\bm{\mu})$. The distance can be chosen flexibly, such as Euclidean distance, cosine similarity. We adopt Euclidean distance, under which the prototypes are equivalent to a Gaussian mixture model in the embedding with an identity covariance matrix (Allen et al., 2019). Meanwhile, temporal pattern $\mathbf{p}$ is used to form the final prediction. Specifically, we fix each prototype temporal pattern $\mathbf{p}$ over a predefined period of length $T$, and during forecasting, the prototype pattern is aligned with the forecast window $H$ based on its phase and length (Lin et al., 2024).